54007GD:

Tess and Dan have $24.00 each to spend at a book fair, where all students receive a 15% discount. They both want to purchase a copy of the same book, which normally sells for $24.50 plus 10% sales tax. To check if she has enough to purchase the book, Tess takes 15% of $24.50 and subtracts that amount from the normal price. She takes 10% of the discounted selling price and adds it back to find the purchase amount. Dan takes 85% of the normal purchase price and then computes 110% of the reduced price. Is Tess correct? Is Dan correct? Do they have enough money to purchase the book? Explain your answer using complete sentences and show your work. Amy, Dan, and Mike are competing in a cooking competition. They all used different amounts of milk from a container holding 6 gallons of milk. • Amy used 13% of the total milk in the container. • Dan used 0.14 of the total milk in the container. • Mike used 0.5 gallon of the total milk in the container. Who used the greatest amount of milk from the container? Show your work and explain your answer in words. The youth group is going on a trip to the state fair. The trip costs $62. Included in that price is $14 for a concert ticket and the cost of 2 passes, one for the rides and one for the game booths. Each of the passes costs the same price. Write an equation representing the cost of the trip, and determine the price of one pass. Solve your equation by showing your work and steps. Andrea has $18 to spend. She spent $6.50, including tax, to buy a notebook. She needs to save $5.75, but she wants to buy a snack. If candy bars cost $1.15 per package including tax, what inequality would show the maximum number of packages that Andrea can buy? Solve your inequality by showing your work and steps.

thatprettygirl:

Tess is wrong while Dan is correct. This is because tax is calculated after discount. Finding the correct amount they are supposed to pay: Price after discount = 85/100 x 24.5 = 20.825 Price after tax = 110/100 x 20.825 = 22.90 Since they have $24 each, it means they can purchase the book.

thatprettygirl:

that was the first one i gave you

54007GD:

i realized afterwards

thatprettygirl:

heres the second one Amy used 13% of the total milk in the container: 0.13*6 = 0.78 gal Dan used 0.14 of the total milk in the container: 0.14*6 = 0.84 gal Mike used 0.5 gallon of the total milk in the container: 0.5 gal 0.5 < 0.78 < 0.84 Dan used the greatest amount of milk from the container.

thatprettygirl:

third one 62=14 + 2x 62-14=2x 48=2x 48/2=x 24 answer: 24

thatprettygirl:

fourth one total cost of the stuff she buys ≤ 18 since she only has 18 dollars to buy stuff let's add up the cost of the stuff she buys: - a flat cost of $6.50 - a savings of $5.75 (she's technically not spending money but we can count this as part of her expenses since it's taking away from the 18 dollars) - some number (x) candy bars. if each candy bar costs $1.15 then the money she spends on candy bars is $1.15